OpenStudy (ria23):

A rumor spreads in a small town with population 20,000 people. The secret spreads throughout the town according to the function, where t represents time, in days: N(t)=20000/10+2490^.97t A.) How many people were in the group that started the rumor? B.) How many people have heard the rumor after 1 day? One week? C.) At this rate, how long will it take for 1950 people to hear the rumor?

OpenStudy (ria23):

\[N(t)=\frac{ 20000 }{ 10+2490e^{-0.97t} }\]

OpenStudy (anonymous):

i should make a template for this one have seen this exact question over a dozen times in the past two years

OpenStudy (anonymous):

A.) How many people were in the group that started the rumor? put \(t=0\) what do you get?

OpenStudy (anonymous):

hint: \(e^0=1\)

OpenStudy (ria23):

It'll just be the exact same problem won't it? 10+2490?

OpenStudy (anonymous):

that is in the denomnator

OpenStudy (anonymous):

\[\frac{20,000}{10+2490}\] compute that one

OpenStudy (coconutjj):

8 people ?

OpenStudy (ria23):

20000/3000 which gives me a long decimal

OpenStudy (anonymous):

\[N(t)=\frac{ 20000 }{ 10+2490e^{-0.97t} }\\ N(0)=\frac{ 20000 }{ 10+2490e^{0} }=\frac{20,000}{10+2490}\]

OpenStudy (anonymous):

oh no dear, what is \(2490+10\) ?

OpenStudy (ria23):

Oh... Woops... Haha 20000/2500 8 c:

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

B.) How many people have heard the rumor after 1 day? \[N(t)=\frac{ 20000 }{ 10+2490e^{-0.97t} }\\ N(1)=\frac{ 20000 }{ 10+2490e^{-.097} }=\text{ use a calculator}\]

OpenStudy (anonymous):

B.) How many people have heard the rumor after 1 week put \(t=7\) compute \[N(7)=\frac{ 20000 }{ 10+2490e^{-.097\times 7} }=\text{ use a calculator}\]

OpenStudy (anonymous):

oops i typed it in wrong in to wolf, let me redo the previous one

OpenStudy (ria23):

The second one is almost 16... 15.714

OpenStudy (anonymous):

same answer though, not quite 9 in any case you can use the wolfram link i sent as a template if you do not want to plug this in to a calculator this would be for 7 days http://www.wolframalpha.com/input/?i=20000%2F%2810%2B2490e^%28-0.097*7%29%29

OpenStudy (anonymous):

yeah that is what i got too

OpenStudy (ria23):

And that's B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

C.) At this rate, how long will it take for 1950 people to hear the rumor? this is the only one that requires any real work any ideas?

OpenStudy (ria23):

would I do the same thing as the last one, but plug in 1950 for t?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

this one says the ANSWER is 1950, what \(t\) did you plug in to get it?

OpenStudy (anonymous):

\[N(t)=\frac{ 20000 }{ 10+2490e^{-0.97t} }=1950\] solve for \(t\)

OpenStudy (ria23):

Can yhu tell me what my first step would be?

OpenStudy (anonymous):

multiply by the denominator to clear the fraction \[20,000=1950(10+e^{-.97t})\]

OpenStudy (anonymous):

damn typo \[20,000=1950(10+2490e^{-97t})\]

OpenStudy (anonymous):

then distribute teh \(1950\) on the right

OpenStudy (anonymous):

or you can divide both sides by 1950 first either way

OpenStudy (ria23):

\[20,000= 19,500+4,855,500e^{-.97t}\]

OpenStudy (ria23):

;-;

OpenStudy (anonymous):

i believe you

OpenStudy (anonymous):

subtract 19500 from both sides

OpenStudy (anonymous):

even i can do that without a calculator \[500=4,855,500e^{-.97t}\]

OpenStudy (ria23):

\[500=4,855,500e^{-.97}\] c:

OpenStudy (anonymous):

divide by \(4,855,500\)

OpenStudy (ria23):

4 million divided by 500? Or 500 divided by the 4 million?

OpenStudy (anonymous):

second way it is a tiny fraction \[\frac{1}{9771}\]

OpenStudy (ria23):

... 9771? I got 9711... ;c

OpenStudy (anonymous):

now comes the only non algebra step \[\frac{1}{9771}=e^{-.97t}\]

OpenStudy (anonymous):

that's cause you did it upside down

OpenStudy (anonymous):

\[a=bx\\ \frac{a}{b}=x\]

OpenStudy (anonymous):

ok now do you know what to do here?\[\frac{1}{9771}=e^{-.97t}\]

OpenStudy (ria23):

No ;-; And I keep getting 9711. ;c I flipped it around and got 1.long decimal ;C

OpenStudy (anonymous):

"no" is a fine answer forget the decimal

OpenStudy (anonymous):

you got \(9771\) because you divided 500 by 4855500

OpenStudy (anonymous):

ack i mean that is what you did not do you divided backwards

OpenStudy (anonymous):

it should be \[500\div 4855500\] and you should get the reciprocal of \(9771\) which is \(\frac{1}{9771}\) a very tiny decimal, lets leave it as a fraction

OpenStudy (ria23):

;c

OpenStudy (ria23):

...........................

OpenStudy (ria23):

Wait the decimal is right?

OpenStudy (anonymous):

oops typo again!

OpenStudy (anonymous):

\[\frac{1}{9711}=e^{-.97t}\\ \ln(\frac{1}{9711})=-.97t\]

OpenStudy (anonymous):

and finally divide by \(-.97\) to find \(t\) i used this http://www.wolframalpha.com/input/?i=ln%281%2F9711%29%2F-.97

OpenStudy (ria23):

Aha! Ok... I'm caught back up. :D So yhu replaced the e with a natural log... And I would divide 1/9711 by -.97?

OpenStudy (anonymous):

take the log of both sides i.e. write in logarithmic form one you have \[\frac{1}{9711}=e^{-.97t}\] you have to get \(t\) out of the exponent you do that via \[\ln(\frac{1}{9711})=-.97t\]

OpenStudy (anonymous):

then simple algebra to get \(t\) namely divide both sides by \(-.97\)

OpenStudy (anonymous):

i sent you the wolfram link with what i am pretty sure is the answer, typo free

OpenStudy (anonymous):

gotta run, good luck

OpenStudy (ria23):

Ok n.n I kind of understand how this was worked out. c: Thank yhu Satellite ^.^